### v1

@@ Abstract
The goal of GBD is to serve as a conduit between data science and empirical research on NP-hard problem classes by facilitating the seamless integration of benchmark data into existing workflows. GBD provides tools for distributing benchmark instances and feature databases, as well as tools for transforming instances and extracting instance features. It also provides a set of prebuilt feature databases for identifying instance equivalence classes, categories, or labels, as well as analytical instance features. Examples of successful applications of GBD include the sanitization and selection of benchmarks for SAT competitions [13], domain-specific solver evaluations [8], and the analysis of solver portfolios and solver prediction models [2]. Moreover, the authors of the latest award-winning state-of-the-art SAT solvers use GBD in their empirical evaluations. [11, 19]. GBD is open source and available under the MIT license [14].
In this paper, we describe the conceptual design of the GBD data model and present the interfaces for interacting with GBD, as well as the extension capabilities provided by the data model, including concrete examples of their use. We start with a brief description of the GBD data model and query language in Section 2. Section 3 presents the GBD interfaces and gives concrete examples of how to use GBD in practice. Finally, Section 4 explains how to extend GBD with additional problem domains, instance formats and feature extractors. We conclude with a summary and an outlook on future work in Section 5.
@@ 2 Data Model and Queries
Context mappings are useful not only for relating instances from different problem domains, but also for relating instances from different contexts within the same problem domain (cf. Use case LABEL:cmd:gbd:sanitize below). Application scenarios include relating different instance formats, analyzing different encodings of application instances, or distinguishing sanitized or otherwise preprocessed instances from their original counterparts.
GBD provides a query language designed to filter instances of a given context by specifying constraints over the instance features. A GBD query is a simple constraint , or a compound constraint , or for queries , , and parentheses can be used to indicate precedence. A simple constraint is of the form with a feature name , a value , and an operator with the usual semantics. For SQL-inspired operators , a constraint can also be of the form , , or to indicate postfix, prefix, or infix constraints, respectively. Value can also be a term, where a term is either a constant number, or a feature name in parentheses 222Parentheses are used here to disambiguate terms and character strings., or a compound term from arithmetic operators and terms , .
@@
Given the cnf context and the data source meta.db providing the instance features track and filename, a query can be of the form to filter for instances from the Anniversary track of the 2022 SAT competition with filenames starting with rphp.
@@ 3.3 Web Interface
The GBD web interface is implemented as a RESTful web service [7] that provides access to the benchmarks and databases of the environment it runs in, and can be started with the command gbd serve. While the GBD command-line tool operates independently of the web service, the web service uses the same GBD API and database configuration as the command-line tool. Our instance of the GBD web interface, accessible via {https://benchmark-database.de}, is configured to run behind an Nginx reverse proxy and is hosted in a Docker container. At the time of writing, we provide access to more than 100 000 benchmark instances from the cnf, wcnf, and opb contexts, with prebuilt feature databases for them.
@@ 3.4 Python Interface
The GBD Python API is wrapped in the class gbd_core.api.GBD and its documentation can be found at https://udopia.github.io/gbd. With the Python API, GBD data sources can be directly integrated into Python scripts and Jupyter notebooks to be used in evaluation and analysis. Use case LABEL:cmd:gbd:api illustrates the use of the Python API to query for instances and features. Line 2 creates a GBD object that connects to the GBD data sources base and meta, which are passed to the constructor as a list of paths. Line 3 creates a list of feature names from the base data source and adds the feature family from the meta data source. Line 4 queries for the given features of cnf instances from the Main track of SAT Competition 2023. The result is returned as a Pandas [16] DataFrame for further analysis.
@@ 4 Writing Extensions
GBD allows for integrating new problem domains by defining new contexts. Contexts are managed in a dictionary data structure that is initialized in the gbd_core.contexts module. The dictionary is indexed with the respective context name and contains a pointer to the instance identification function and a list of valid instance filename extensions. The instance identifier function is the function that assigns a unique instance identifier to each benchmark instance, which serves as the primary key in the context-specific feature databases. The list of valid instance filename extensions is used to locate benchmark instances in the local file system during the database initialization process. Once a context is created in the dictionary, context-specific databases can be created and populated using the context name.
New feature extractors can be integrated into GBD by defining feature extractor functions. Feature extractor functions are managed in a dictionary data structure that is initialized in the gbd_init.feature_extractors module. The dictionary is indexed by the name of the feature extractor and contains a list of tuples of feature names and default values, a list of valid contexts in which it can be used, and a pointer to the feature extractor function. All registered feature extractors are accessible by their names as subcommands of gbd init.
Another way to introduce new instances and features to GBD is to implement an instance transformer. Instance transformers are managed in a dictionary data structure that is initialized in the gbd_init.instance_transformers module. The dictionary is indexed by the name of the instance transformer and contains a list of valid source and target contexts, a list of tuples of the names of the created features and their default values, a pointer to a function to get the name of the instance file to create, and a pointer to the instance transformer function itself. An instance transformer function takes an instance in one context and transforms it to an instance in another context. The new instance is hereby simultaneously integrated in the database and additional features can be stored.

### v2

@@ Abstract
In essence, the purpose of GBD is to act as a conduit between data science and empirical research on NP-hard problem classes by facilitating the seamless integration of benchmark data into existing workflows. GBD provides tools for distributing benchmark instances and feature databases, as well as tools for transforming instances and extracting instance features. It also provides a set of prebuilt feature databases for identifying instance equivalence classes, categories, or labels, as well as analytical instance features. Examples of successful applications of GBD include the sanitization and selection of benchmarks for SAT competitions [13], domain-specific solver evaluations [8], and the analysis of solver portfolios and solver prediction models [2]. Moreover, the authors of the latest award-winning state-of-the-art SAT solvers use GBD in their empirical evaluations. [11, 21]. GBD is open source and available under the MIT license [14].
In this paper, we describe the conceptual design of the GBD data model and present the interfaces for interacting with GBD, as well as the extension capabilities provided by the data model, including concrete examples of their use. We start with a brief description of the GBD data model and query language in Section 2. Section 3 presents the GBD interfaces and gives concrete examples of how to use GBD in practice. Finally, Section 4 explains how to extend GBD with additional problem domains, instance formats, and feature extractors. We conclude with a summary and an outlook on future work in Section 5.
@@ 2 Data Model and Queries
Context mappings are useful not only for relating instances from different problem domains, but also for relating instances from different contexts within the same problem domain. Application scenarios include relating different instance formats, analyzing different encodings of application instances, or distinguishing sanitized or otherwise preprocessed instances from their original counterparts (cf. Use case LABEL:cmd:gbd:sanitize below).
Queries.
GBD provides a query language designed to filter instances of a given context by specifying constraints over the instance features. A GBD query is a simple constraint , or a compound constraint , or for sub-constraints , , and parentheses can be used to indicate precedence. A simple constraint is of the form with a feature name , a value , and an operator with the usual semantics. For SQL-inspired operators , a constraint can also be of the form , , or to indicate postfix, prefix, or infix constraints, respectively. Value can also be a term, where a term is either a constant number, or a feature name in parentheses 222Parentheses are used here to disambiguate terms and character strings., or a compound term from arithmetic operators and terms , .
@@
Given the cnf context and the data source meta.db providing the instance features track and filename, a query can be of the form to filter for instances from the Anniversary track of the 2022 SAT competition whose filenames begin with the string rphp.
@@ 3.3 Web Interface
The GBD web interface is implemented as a RESTful web service [7] that provides access to the benchmarks and databases of the environment it runs in, and can be started with the command gbd serve. While the GBD command-line tool operates independently of the web service, the web service uses the same GBD API and database configuration as the command-line tool. Our instance of the GBD web interface, accessible via {https://benchmark-database.de}, is configured to run behind an Nginx reverse proxy [20] and is hosted in a Docker container [17]. At the time of writing, we provide access to more than 100 000 benchmark instances from the cnf, wcnf, and opb contexts, with prebuilt feature databases for them.
@@ 3.4 Python Interface
The GBD Python API is wrapped in the class gbd_core.api.GBD and its documentation can be found at https://udopia.github.io/gbd. With the Python API, GBD data sources can be directly integrated into Python scripts and Jupyter notebooks to be used in evaluation and analysis. Use case LABEL:cmd:gbd:api illustrates the use of the Python API to query for instances and features. Line 2 creates a GBD object that connects to the GBD data sources base and meta, which are passed to the constructor as a list of paths. Line 3 creates a list of feature names from the base data source and adds the feature family from the meta data source. Line 4 queries for the given features of cnf instances from the Main track of SAT Competition 2023. The result is returned as a Pandas [16] DataFrame for further analysis.555We provide illustrative examples of benchmark data analyses, including portfolio analysis, category prediction, and category-specific ranking, accessible via https://udopia.github.io/gbdeval/.
@@ 4 Writing Extensions
The architectural design of GBD allows the integration of new contexts, feature extractors, and instance transformers. The extensibility of the system is achieved through the use of dictionaries, which serve as registries for contexts, feature extractors, and instance transformers. The following section describes the elements required to create such a registry entry. It is important to note that at the time of writing, these registries are hard-coded, which is a limitation in that it requires modification of these dictionaries within the source code of GBD. This issue will be addressed in future versions of GBD, where the registry will be moved to configuration files.
GBD contexts are managed in a dictionary data structure that is initialized in the gbd_core.contexts module. The dictionary is indexed with the respective context name and contains a pointer to the instance identification function and a list of valid instance filename extensions. The instance identification function is the function that assigns a unique instance identifier to each benchmark instance, which serves as the primary key in the context-specific feature databases. The list of valid instance filename extensions is used to locate benchmark instances in the local file system during the database initialization process. Once a context is created in the dictionary, context-specific databases can be populated using the context name.
A new feature extractor can be integrated into GBD by registering a feature extractor function in a dictionary data structure that is initialized in the gbd_init.feature_extractors module. The dictionary is indexed by the name of the feature extractor and contains a list of the provided features, which are tuples of feature names and default values. It also contains a pointer to the feature extractor function and a list of contexts in which it can be used. Registered feature extractors are automatically accessible by their names as subcommands of gbd init.
Instance transformers are managed in a dictionary data structure that is initialized in the gbd_init.instance_transformers module. An instance transformer function takes an instance in a source context and transforms it to a new instance in the target context. The dictionary is indexed by the name of the instance transformer and, similarly to feature extractors, contains a list of the provided features. Additionally, the dictionary contains a pointer to the instance transformer function itself, as well as a pointer to a function that generates the name of the benchmark instance to be created. Registered instance transformers are automatically accessible by their names as subcommands of gbd transform.